# Astrometry-Only Detection of Microlensing Events with Gaia

### v1

@@ Acknowledgements.
T. J., A. G., Ł. W., T. P., M. K., and M. L. acknowledge the financial support from European Space Agency PRODEX Experiment Arrangement GAME. T. J., A. G., T. P., M. B., and M. K. acknowledge the financial support from the Slovenian Research Agency (research core funding P1-0031, infrastructure program I0-0033, and project grants Nos. J1-8136, J1-2460, N1-0344). Ł. W. acknowledges the funding from the European Union’s research and innovation programme under grant agreements Nos. 101004719 (ORP) and 101131928 (ACME). We acknowledge ESA Gaia and the many researchers from Gaia DPAC for their valuable contributions and expertise at different phases of this project. Researcher T. J. conducts his research under the Marie Skłodowska-Curie Actions – COFUND project, which is co-funded by the European Union (Physics for Future – Grant Agreement No. 101081515). Z. K. is a Fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). The following software was used in this work: astromet (https://github.com/zpenoyre/astromet.py), jaxtromet (https://github.com/maja-jablonska/jaxtromet.py), jaxopt (Blondel et al. 2021), Matplotlib (Hunter 2007).
@@ Appendix A GAME Filter
In the following, we describe a method for obtaining initial guesses for individual parameters $\alpha_{0}^{0}$, $\delta_{0}^{0}$, $\mu_{\alpha^{\star}}^{0}$, $\mu_{\delta}^{0}$, $\varpi^{0}$, $u_{0}^{0}$, $\theta_{\mathrm{E}}^{0}$, $t_{0}^{0}$, $t_{\mathrm{E}}^{0}$, $\pi_{\mathrm{EE}}^{0}$, and $\pi_{\mathrm{EN}}^{0}$, and the main steps of the minimization process.
@@ A.2 Minimization steps
    Calculate $L_{\mathrm{opt}}$ after minimization.
@@
Figure 13: Distribution of $L_{\mathrm{opt}}$ for the lens_G14_N281_DR4 dataset (see Tables 1 and 3). Blue and green colours denote events with $L_{\mathrm{opt}}\leq 0.01$ and $L_{\mathrm{opt}}>0.01$, respectively. The values in individual bins denote the number of events within the corresponding bin range.

### v2

@@ Acknowledgements.
T. J., A. G., Ł. W., T. P., M. K., and M. L. acknowledge the financial support from the European Space Agency PRODEX Experiment Arrangement GAME. T. J., A. G., T. P., M. B., and M. K. acknowledge the financial support from the Slovenian Research Agency (research core funding P1-0031, infrastructure program I0-0033, and project grants Nos. J1-8136, J1-2460, N1-0344). Ł. W. acknowledges the funding from the European Union’s research and innovation programme under grant agreements Nos. 101004719 (ORP) and 101131928 (ACME). We acknowledge ESA Gaia and the many researchers from Gaia DPAC for their valuable contributions and expertise at different phases of this project. Researcher T. J. conducts his research under the Marie Skłodowska-Curie Actions – COFUND project, which is co-funded by the European Union (Physics for Future – Grant Agreement No. 101081515). Z. K. is a Fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). The following software was used in this work: astromet (https://github.com/zpenoyre/astromet.py), jaxtromet (https://github.com/maja-jablonska/jaxtromet.py), jaxopt (Blondel et al. 2021), Matplotlib (Hunter 2007).
@@ Appendix A GAME Filter
In the following, we describe a method for obtaining initial guesses for individual parameters $\alpha_{0}^{0}$, $\delta_{0}^{0}$, $\mu_{\alpha^{\star}}^{0}$, $\mu_{\delta}^{0}$, $\varpi^{0}$, $u_{0}^{0}$, $\theta_{\mathrm{E}}^{0}$, $t_{0}^{0}$, $t_{\mathrm{E}}^{0}$, $\pi_{\mathrm{EE}}^{0}$, and $\pi_{\mathrm{EN}}^{0}$, and the main steps of the minimisation process.
@@ A.2 Minimisation steps
    Calculate $L_{\mathrm{opt}}$ after minimisation.
@@
Figure 13: Distribution of $L_{\mathrm{opt}}$ for the lens_G14_N281_DR4 mock dataset (see Tables 1 and 3). Blue and green colours denote events with $L_{\mathrm{opt}}\leq 0.01$ and $L_{\mathrm{opt}}>0.01$, respectively. The values in individual bins denote the number of events within the corresponding bin range.